vernier calipers viva

Grade:12th Pass

1 Answers

Komal
8 years ago
Q.1: What is vernier calipers?
Ans. It is simple instrument by which length of an object can be measured accurately up to one-tenth of a millimeter or one-hundredth of a centimeter.

Q.2: What is the use of Vernier Scale?
Ans. To increase the accuracy of the main scale.

Q.3: What is least count (L.C.)?
Ans. The minimum measurement that can be taken by an instrument.

Q.4: What is the least count of?
(a) meter scale (b) Stop Watch (c) Vernier Calliper
Ans. The least count of:
(a) meter scale = 0.1 cm or 1 mm.
(b) stop watch = 0.01 sec or 1/100 sec. or according to stop watch.
(c) vernier scale = 0.01 cm or 0.1 mm

Q.5: What are units of vernier scale?
Ans. Vernier has not units of its own.

Q.6: What is the difference in the measurement of one smaller division on the main scale (M.S.) and that of vernier scale (V.S.)
Ans. (1 smallest Div. on M.S.) – (1 smallest Div. on V.S) = 0.1 – 0.09 = 0.01 cm = Least count.

Q.7: How the L.C. of a vernier is determined?
Ans. By using the relation:
L.C. = (Smallest division on Main Scale ) / (Total number of divisions on Vernier Scale)

Q.8: Will the accuracy of a vernier will increase or decrease if 20 divisions are made in the same length?
Ans. The accuracy will increase since the L.C. will decrease i.e. Accuracy α = 1 / Least count

Q.9: Which one is smaller M.S. div. or V.S. div.?
Ans. V.S. division.

Q.10: 1 division on V.S. = _____ cm.
Ans. 0.09 cm.
10 div. on V.S. = 0.9 cm
1 div. on V.S. = 0.9/10 cm = 0.09 cm

Q.11: What are the functions of (a) Sliding strip (b) Upper jaws
Ans. (a) For measuring depth. (b) For measuring the inner diameter.

Q.12: What is meant by zero error (Z.E.)?
Ans. The error which arises when zeros of M.S. and V.S. do not coincide upon joining the two jaws.

Q.13: When Z.E. is positive?
Ans. The Z.E. is positive when the zero of V.S. is to the right of the zero of M.S.

Q.14: How we find the corrected reading?
Ans. Corrected reading = Meaning reading – ( + Z.E.)

Q.15: Do you know any use of vernier scale?
Ans. It is used in spectrometer, barometer, traveling microscope, etc.

Q.16: What are significant figures?
Ans. Numbers 0, 1, 2, 3, ……. are called significant figures when they occur. The zero is not a significant figure if it is at the extreme left or extreme right of a number, e.g. 1.02 has three significant figures, but 1.20 has two significant figures.

Q.17: What is meant by degree of accuracy?
Ans. It is the measurement estimated by the number of significant figures in the result.

Q.18: Why is the vernier calipers so called?
Ans. A French mathematician named Pierre Vernier invented it and it is named after him.

Q.19: How does temperature affect the length of a cylinder?
Ans. The length of a cylinder increases with the increase of temperature and decreases with the
decrease of temperature.

Q.20: What is the unit of volume?
Ans. In MKS system, the unit of volume is m3. In CGS system, the unit of volume is cm3. In FPS
system, the unit of volume is ft3.